(1 point) For the following system to be consistent, –5x– 7y=5z = 5 17x-49y+kz =-10 –8x+ 7y+5z = 6 we must have, k =
Answer #1

«»;; ;; ܕܕܐ ܙ; ܢ ܟܐ -5x - 74 - 5z = 5 17*-499 +K7 = -10 -82 +7y +52 = 6 we have to find ka? for which the above system in ma1-5 -7 - 5 -36476 K-17 7 TO +35 ! ! So for this to be consistent we must have K²-35 Otherwise of K=-35 we get 1-5 -7. -s 157